The blood groups of 200 people are distributed as follows: 50 have type A blood, 65 have B blood type, 70 have O blood type and 15 have type AB blood. If a person from this group is selected at random, what is the probability that this person has O blood type?

Construct a table of frequencies for the blood groups


1
Expert's answer
2021-11-15T20:28:43-0500

Table of frequencies is as given below

Blood group Frequency

A 50

B 65

O 70

AB 15


Total frequency is 50+65+70+15=200

If a person is randomly selected from this group, the probability that he has O blood type is,

sample size is 200

Number of people with blood group O is 70

Probability that the randomly selected person has blood type O is 70/200=7/20.

Thus, the probability that this person has O blood type is 7/20.